OpenStudy (anonymous):

The area of a square is (9x2 - 12x + 4) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work

OpenStudy (fibonaccichick666):

in order to get a \(9x^2\) you have to have (____+b)(_____+a)

OpenStudy (fibonaccichick666):

the blanks are the same

OpenStudy (anonymous):

The same as what?

OpenStudy (fibonaccichick666):

the same value

OpenStudy (anonymous):

I still don't understand

OpenStudy (anonymous):

would we do the same exact thing that we did in the previous equation?

OpenStudy (fibonaccichick666):

yea sort of that example is similar

OpenStudy (fibonaccichick666):

you first want to break the 9x^2 into two parts

OpenStudy (fibonaccichick666):

because whatever \(\sqrt{9x^2}\) is, that will fill in the blanks

OpenStudy (fibonaccichick666):

this might be more like the third problem

OpenStudy (anonymous):

Could we split it in 3x and 3x?

OpenStudy (fibonaccichick666):

yup

OpenStudy (fibonaccichick666):

so now you need the a and b

OpenStudy (anonymous):

so, (3x + b)(3x + a)

OpenStudy (fibonaccichick666):

yea exactly

OpenStudy (fibonaccichick666):

so now, you know that ab=4 and that 3a+3b=-12

OpenStudy (fibonaccichick666):

so you want to solve the system of eq for the answer

OpenStudy (fibonaccichick666):

hint, you only have one option really

OpenStudy (anonymous):

-2

OpenStudy (fibonaccichick666):

and ?

OpenStudy (fibonaccichick666):

so our final answer is?

OpenStudy (anonymous):

3(-2) + 3(-2) = -12 a = -2 b = -2

OpenStudy (fibonaccichick666):

good, so now what is the factored poly?

OpenStudy (anonymous):

I don't know what you are asking for?

OpenStudy (fibonaccichick666):

what two things do you multiply together to get that original polynomial(poly for short)?

OpenStudy (anonymous):

(3x - 2)^2

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

Thank you! and sorry about the late replies, my connection keeps getting lost.

OpenStudy (fibonaccichick666):

it's cool, np